# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

### v1

@@ 2.1 Challenges in Visual Reasoning: Limitations of LVLMs
Method Model
Llama3-LLaVA-NeXT-8B LLaVA-OneVision-Qwen2-7B-OV Qwen-VL-Chat
Direct 41.8 49.1 36.2
CoT 41.5 46.5 35.4
VDGD 42.3 49.7 37.1
@@
2.2 Drawbacks of Passive Information Extraction
For further demonstration, we analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information necessary for reasoning, respectively. Similar to Sec. 2.1, we pair the image captions with the reasoning process of GPT-4o-mini, and employ GPT-4 as a judge to rate the captions from 1 to 5 across these dimensions. The prompt template is shown in Figure 5. As shown in Table 3, the captions for correct responses of Llama3-LLaVA-NeXT-8B receive higher scores across all three criteria, highlighting the importance of better captions for multi-modal reasoning. Additionally, all captions score significantly lower in the Question Relevance and Reasoning Effective Info Inclusion dimensions than the Detail Level dimension, indicating that while the captions are detailed, they often lack relevance to the questions. Figure 2 shows a case where Llama3-LLaVA-NeXT-8B utilizes fine-grained captions to solve a question from the MMMU benchmark. As illustrated, although the caption exhaustively describes the image content, it incorrectly describes the wires in the image as octagons, and misses information about the locations of these wires. This information is irrelevant to the target question, thus offering minimal assistance to LVLMs. In summary, our analysis highlights the drawbacks of passive visual reasoning enhancement techniques in terms of information insufficiency and redundancy, due to their question-agnostic property.
(1) ProReason
(2) Passive Visual Reasoning Enhancement Technique
(3) Multi-step Reasoning Framework for LLMs
Figure 1: Overview and comparison of workflows of ProReason, VDGD and ReAct. Unlike existing works (e.g., VDGD and ReAct), our proposed method enables the model to actively acquire necessary information from the images, making it more suitable for LVLMs.
3 Method
As illustrated in Figure 1, ProReason consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component. The entire workflow comprises three steps: Action, Judgment, and Summary, effectively integrating the proactive visual understanding (i.e., eyesight) and textual reasoning (i.e., wisdom) capacities for enhanced multi-modal reasoning performance.
3.1 Action Step
Action step is the core of question-oriented visual information extraction, driven by three sub-agents: Dispatcher, Vision Expert, and Reasoning Expert. The Dispatcher orchestrates the workflow, selectively directing the Vision Expert to capture specific visual information, or instructing the Reasoning Expert to analyze known information to derive more. The responses from both the Vision Expert and the Reasoning Expert are stored in a textual Memory component.
Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Reasoning Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty). The Dispatcher then generates a query $q$ for the chosen expert. If the Vision Expert is selected, it takes the image $I$ and query $q$ as input, and generates an answer $A_{v}$, which is then stored in the Memory. When the Reasoning Expert is selected, it provides a response consisting of the reasoning process and the final answer $A_{r}$ based on the query $q$ and known information in the Memory, before only $A_{r}$ is stored in Memory. Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT and ReAct, thereby suffering less from redundant information.
3.2 Judgment Step
Judgment step is conducted by a sub-agent called Referee, which evaluates whether the information stored in Memory is sufficient to answer the question $Q$. The input to the Referee includes the question $Q$ and available information in the Memory. If the Memory contains adequate information to answer the question $Q$, the Referee outputs the identifier “SOLVABLE”; otherwise, it outputs “UNSOLVABLE”. If the Referee’s output is SOLVABLE, the workflow precedes to the Summary phase. Conversely, if the output is UNSOLVABLE, the Action phase is re-executed to gather more necessary information. This allows the Referee to collaborate closely with the three sub-agents in the Action step, enabling the framework to proactively acquire the necessary information and prevent omissions, thereby overcoming the drawbacks of passive methods.
3.3 Summary Step
The Summary step focuses on integrating the available information in the Memory, and providing the final answer to the question $Q$. This step is mainly powered by a sub-agent called Summarizer. Once the Referee determines that the information in the Memory is sufficient to address the question $Q$, and outputs the identifier SOLVABLE, the Summarizer will be called to draw a conclusion, based on the Memory information. This conclusion represents the answer to the question $Q$ of the whole framework, and will be evaluated by the performance metrics.
3.4 Advantages of ProReason
Reduced information Mission or Redundancy.
With the close collaboration of the Dispatcher, Vision Expert, Reasoning Expert, and Referee agents, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy. Meanwhile, the Memory component retains only the image captions from the Vision Expert and the reasoning results from the Reasoning Expert, providing a compact textual descriptions. This contributes to minimize the interference of irrelevant information on the subsequent Summarizer.
LLM-assisted Multi-modal Reasoning.
In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents. These agents are then effectively organized through a designate pipeline to complete the tasks. This decomposition-before-integration approach not only circumvents the inherent differences between modalities, but also allows visually irrelevant sub-agents to be performed by text-only LLMs. Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power multi-modal reasoning process for performance enhancement.
4 Experiments
4.1 General Setup
Datasets.
To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a). All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly as follows:
  • MME is an inclusive benchmark that encompasses 14 subtasks, designed to evaluate perceptual and cognitive abilities. Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning. [^2] To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers. For the results using the official MME calculation method, please refer to Table 9, Table 10, Table 11, and Table 12.
Base Models.
We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources. As one of the most performant LVLMs, GPT-4o-mini demonstrates significant advancements in visual reasoning capabilities, and provides cheap and fast API. In contrast, Llama3-LLaVA-NeXT-8B is a fully open-sourced LVLM developed by the LLaVA team based on the Llama-3-8B LLM (Dubey et al., 2024) and CLIP vision encoder (Radford et al., 2021). Qwen2.5-72B-Instruct is a robust LLM designed to deliver high-quality instruction following and handle complex tasks. It will be utilized for achieving LLM-assisted multi-modal reasoning.
Baselines.
Implementation Details.
The prompt templates for all methods are shown in Figures 6, 7 and 8. Specifically for ProReason, in order to prevent an infinite loop, if Dispatcher selects Vision Expert or Reasoning Expert to obtain information up to 5 times, and Referee still judges that the existing knowledge in Memory is not enough to solve the question, the information collection phase will be terminated immediately, and then Summarizer will directly answer the question based on the available information in Memory. This setup, refined through multiple trials, is the most effective. Additionally, since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini.
Model Method Dataset
MME MMMU MathVista Hallu.
Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6)
CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4)
CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7)
ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2)
ProReason 71.5 (+10.0) 52.5 (+10.7) 38.8 (+1.7) 50.9 (+5.1)
Average 66.30 44.22 36.06 44.16
GPT-4o-mini Direct 79.2 48.4 53.0 56.0
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9)
ProReason 90.4 (+11.2) 61.6 (+13.2) 54.9 (+1.9) 58.9 (+2.9)
Average 84.63 54.82 52.63 55.23
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
ReAct 68.5 46.7 31.7 43.6
ProReason 71.5 52.5 38.8 50.9
ProReason 81.3 (+9.8) 56.8 (+4.3) 48.8 (+10.0) 52.3 (+1.4)
GPT-4o-mini ReAct 73.6 (+5.1) 48.4 (+1.7) 36.2 (+4.5) 46.7 (+3.1)
ProReason 84.7 (+13.2) 54.5 (+2.0) 41.7 (+2.9) 53.1 (+2.2)
4.2 Main Results
4.3 Relative Importance of Sub-agents
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Dispatcher Vision Expert Reasoning Expert Referee Summarizer
MME 90.4 88.8 (-1.6) 84.7 (-5.7) 88.7 (-1.7) 89.6 (-1.1) 84.2 (-6.2)
MMMU 61.6 60.9 (-0.7) 54.5 (-7.1) 60.2 (-1.4) 51.5 (-10.1) 51.0 (-10.6)
Summarizer is the most crucial sub-agent, closely followed by Referee. The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively. This highlights the critical function of the Summarizer in integrating all available information to conclude final answers. Besides, the substitution of Referee leads to a 10.1% reduction on MMMU. Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
MME 84.7 (+13.2) 77.8 (+6.3) 71.5 58.8 61.5
MMMU 54.5 (+2.0) 53.4 (+0.9) 52.5 41.5 41.8
4.4 Which One is More Crucial: Visual Understanding or Textual Reasoning?
5 Related work
Large Visual-Language Model.
Recently, large vision-language models (LVLMs) (Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents. In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c; b; 2024a; Li et al., 2024a; b) and InternVL (Chen et al., 2024b) families, have been extensively developed. In the closed-source domain, proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini Pro 1.5 (Reid et al., 2024) have also achieved significantly success. Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously. This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Ghosh et al., 2024) and motivating more effective solutions.
Passive Visual Reasoning.
Multi-step Reasoning Framework.
6 Conclusion
In this paper, we first validate that existing multi-modal reasoning approaches still suffer insufficient and irrelevant visual descriptions, as well as limited multi-modal capacities. To address these issues, we decompose the visual reasoning process into visual perception and textual reasoning stages, and introduce a novel visual reasoning framework named ProReason, featuring multi-run proactive perception and decoupled vision-reasoning capabilities. Empirically, extensive experiments demonstrate the superiority of ProReason over both passive image information acquisition methods and multi-step reasoning frameworks for text-only LLMs across multiple visual reasoning benchmarks with both open-source and closed-source models. Notably, our method showcases the remarkable feasibility of integrating LLMs for multi-modal reasoning with dramatically improved performance, highlighting the great potential for LLM-assisted LVLM reasoning in future research.
Appendix A Appendix
A.1 Demonstrative Examples
A.2 Prompts
Figure 7: Prompt templates of Dispatcher.
A.3 Supplementary Results
@@
Llama3-LLaVA- NeXT-8B Direct 359 41.8 37.1 45.8
VDGD 374 42.3 36.1 44.2
CCoT 354 40.5 36.8 37.4
CoT 258 41.5 35.9 43.1
ReAct 385 46.7 31.7 43.6
ProReason 416 52.5 38.8 50.9
GPT-4o-mini Direct 552 48.4 53.0 56.0
VDGD 584 51.4 51.2 52.4
CCoT 535 54.2 53.6 56.7
CoT 729 58.5 53.8 56.3
ReAct 712 54.8 49.3 51.1
ProReason 779 61.6 54.9 58.9
@@
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Dispatcher Vision Expert Reasoning Expert Referee Summarizer
MME 779 746 613 743 762 641
MMMU 61.6 60.9 54.5 60.2 51.5 51.0

### v2

@@ A.3.5 Frequency of selection of various experts
Model Method
Direct CoT VDGD
Llama3-LLaVA-NeXT-8B 41.8 41.5 42.3
LLaVA-OneVision-Qwen2-7B-OV 49.1 46.5 49.7
Qwen-VL-Chat 36.2 35.4 37.1
@@
Model Agent Dataset
MME MMMU
GPT-4o-mini 90.4 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
Vision Expert 84.7 (-5.7) 54.5 (-7.1)
Reasoning Expert 88.7 (-1.7) 60.2 (-1.4)
Referee 89.6 (-1.1) 51.5 (-10.1)
Summarizer 84.2 (-6.2) 51.0 (-10.6)
Table 13: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments. The blue text highlights the performance degradation due to sub-agent integration.
GPT-4o-mini Dataset
MMMU MME
ProReason 61.6 91.9
CoT 58.5 87.8
Vision & Reasoning
Expert Integration
58.6 (-3.0) 90.4 (-1.5)
Dispatcher,
Vision & Reasoning Expert,
& Referee Integration
57.8 (-3.8) 90.8 (-1.1)
All Sub-Agents Integration 56.8 (-4.8) 89.2 (-2.7)
@@
Llama3-LLaVA- NeXT-8B Direct 35.9 41.8 37.1 45.8
VDGD 37.4 42.3 36.1 44.2
CCoT 35.4 40.5 36.8 37.4
CoT 25.8 41.5 35.9 43.1
ReAct 38.5 46.7 31.7 43.6
ProReason 41.6 50.5 38.8 50.9
GPT-4o-mini Direct 55.2 48.4 53.0 56.0
VDGD 58.4 51.4 51.2 52.4
CCoT 53.5 54.2 53.6 56.7
CoT 72.9 58.5 53.8 56.3
ReAct 71.2 54.8 49.3 51.1
ProReason 77.9 61.6 54.9 59.9
@@
Model Agent Dataset
MME MMMU
GPT-4o-mini 779 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 746 60.9
Vision Expert 613 54.5
Reasoning Expert 743 60.2
Referee 762 51.5
Summarizer 641 51.0
Table 18: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments. Different from Table 13, MME scores are calculated with the official method here.
GPT-4o-mini Dataset
MMMU MME
ProReason 61.6 779
CoT 58.5 729
Vision & Reasoning
Expert Integration
58.6 673
Dispatcher,
Vision & Reasoning Expert,
& Referee Integration
57.8 714
All Sub-Agents Integration 56.8 711
A.4 Demonstrative Examples
Figure 3: A complete reasoning process of ProReason for the case shown in Figure 2.
Figure 6: A typical mistake made by ProReason. The vision expert incorrectly identifies 4:30 as 6:25, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
Figure 7: A typical mistake made by ProReason. The vision expert incorrectly provided the coordinates of the three vertices of the triangle, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
A.5 Prompts
Figure 10: Prompt templates of Dispatcher.
Figure 12: Prompt templates of LLMs-based ProReason evaluation.
Figure 13: Prompt templates of Vision and Reasoning Expert Integration.
Figure 14: Prompt templates of Dispatcher, Vision Expert Reasoning Expert and Referee Integration.
Figure 15: Prompt templates of All Five Sub-Agent Integration.